# Learning to Kern — Set-wise Estimation of Optimal Letter Space

### v1

@@ Abstract
Kerning is the task of setting appropriate horizontal spaces for all possible letter pairs of a certain font. One of the difficulties of kerning is that the appropriate space differs for each letter pair. Therefore, for a total of 52 capital and small letters, we need to adjust $52\times 52=2704$ different spaces. Another difficulty is that there is neither a general procedure nor criterion for automatic kerning; therefore, kerning is still done manually or with heuristics. In this paper, we tackle kerning by proposing two machine-learning models, called pairwise and set-wise models. The former is a simple deep neural network that estimates the letter space for two given letter images. In contrast, the latter is a Transformer-based model and estimates the letter spaces for three or more given letter images. For example, the set-wise model simultaneously estimates 2704 spaces for 52 letter images for a certain font. Among the two models, the set-wise model is not only more efficient but also more accurate because its internal self-attention mechanism allows for more consistent kerning for all letters. Experimental results on about 2500 Google fonts and their quantitative and qualitative analyses show that the set-wise model has an average estimation error of only about 5.3 pixels when the average letter space of all fonts and letter pairs is about 115 pixels.
@@ 1 Introduction
Kerning, or letter spacing, is the micro-typographic task of adjusting the horizontal space between adjacent letters. For monospaced fonts (like most Chinese fonts), the space is easily determined so that all letters occupy the same horizontal space. In contrast, kerning is a very important operation for the proportional fonts of the Latin alphabet letters. Fig. 1(a) shows the effect of kerning. Without kerning (that is, by monospacing), the letters do not appear to form the word “UNIVERSITY,” especially due to an excessive space around ‘I.’ Various cognitive psychology research on reading has revealed that kerning is important for better readability [1, 4]. Therefore, kerning is not a trivial task for type designers and is beneficial for readers.
@@
Currently, kerning is determined manually or by heuristic rules. For the manual cases, kerning is an enormous task; for the 52 letters of the Latin alphabet (i.e., ‘A’ to ‘z’), type designers need to determine the space for each of $270452^{2}=$ 2704 possible letter pairs. To relax this task, heuristic rules have also been used. A well-known heuristic is optical spacing [7], which tries to keep the blank space area between adjacent letters constant. (For example, the blank space between ‘A’ and ‘V’ in Fig. 1(b) is the parallelogram-like region between them. Its area is finally measured with additional perceptual heuristics [9] [^1] A practical implementation of optical spacing is the “hz-program” [9, 16, 19] by a German type foundry called URW. Unfortunately, URW has not published the details of the hz-program [16]..) It should be noted that heuristic rules are imperfect — after applying the rules, a final adjustment process by type designers is still required. For example, although FontForge is a de-facto standard font design software and has a heuristics-based automatic kerning mode, users need to revise its kerning result.
@@ 4.3 Comparative methods and evaluation metrics
  • “Average” determines the letter space between a certain letter pair (say, “AB”) by averaging the ground-truth spaces of that pair (“AB”) from all fonts in the training set. Consequently, the pair “AB” always has the same space regardless of font, whereas the spaces for “AB” and “IJ” are different.
@@ 4.5 Qualitative evaluation
These results, especially those with small MAEs (less than about two pixels), show that the set-wise model performs near-perfect space estimation. More importantly, the results are not tragic even when larger MAEs. In fact, in Fig. 9, the word “UNIVERSITY” with the estimated spaces has no extra space to split the word into fragments and no under-estimated space to overlap adjacent letters, except for Mrs Sheppards Regular, where letters in a Handwriting style are overlapping. Another interesting result is Lekton Italic, which becomes the worst due to the under-estimated spaces of several pairs; however, the under-estimated spaces still seem also (or even more) natural.
@@ 6 Conclusion, Limitation, and Future Work
Our current limitation is the inflexibility of the estimation output. In this paper, we treated the default letter spaces as the ground-truth and tried to estimate the letter spaces similar to the ground-truth. In other words, we assumed that the appropriate letter space is uniquely determined for each letter pair of each font. However, typographic designers modify the letter spaces to create a particular impression. For example, letter spacing slightly wider than the default gives a more established and confident impression. In future work, such a space modification mechanism should be introduced into the current model to allow for more flexible kerning. We also plan to apply our model to font images generated by GANs or diffusion models.
Acknowledgment: This work was supported by JSPS KAKENHI Grant Number JP22H00540.

### v2

# Learning to Kern: Set-wise Estimation of Optimal Letter Space

@@ Abstract
Kerning is the task of setting appropriate horizontal spaces for all possible letter pairs of a certain font. One of the difficulties of kerning is that the appropriate space differs for each letter pair. Therefore, for a total of 52 capital and small letters, we need to adjust $52\times 52=2704$ different spaces. Another difficulty is that there is neither a general procedure nor criterion for automatic kerning; therefore, kerning is still done manually or with heuristics. In this paper, we tackle kerning by proposing two machine-learning models, called pairwise and set-wise models. The former is a simple deep neural network that estimates the letter space for two given letter images. In contrast, the latter is a transformer-based model that estimates the letter spaces for three or more given letter images. For example, the set-wise model simultaneously estimates 2704 spaces for 52 letter images for a certain font. Among the two models, the set-wise model is not only more efficient but also more accurate because its internal self-attention mechanism allows for more consistent kerning for all letters. Experimental results on about 2500 Google fonts and their quantitative and qualitative analyses show that the set-wise model has an average estimation error of only about 5.3 pixels when the average letter space of all fonts and letter pairs is about 115 pixels.
@@ 1 Introduction
Kerning, or letter spacing, is the micro-typographic task of adjusting the horizontal space between adjacent letters[10]. For monospaced fonts (like most Chinese fonts), the space is easily determined so that all letters occupy the same horizontal space. In contrast, kerning is a very important operation for the proportional fonts of the Latin alphabet letters. Fig. 1(a) shows the effect of kerning. Without kerning (that is, by monospacing), the letters do not appear to form the word “UNIVERSITY,” especially due to an excessive space around ‘I.’ Various cognitive psychology research on reading has revealed that kerning is important for better readability [1, 4]. Therefore, kerning is not a trivial task for type designers and is beneficial for readers.
@@
Currently, kerning is determined manually or by heuristic rules. For the manual cases, kerning is an enormous task; for the 52 letters of the Latin alphabet (i.e., ‘A’ to ‘z’), type designers need to determine the space for each of $270452^{2}=$ 2704 possible letter pairs. To relax this task, heuristic rules have also been used. A well-known heuristic is optical spacing [7], which tries to keep the blank space area between adjacent letters constant. (For example, the blank space between ‘A’ and ‘V’ in Fig. 1(b) is the parallelogram-like region between them. Its area is finally measured with additional perceptual heuristics [9] [^1] The “hz-program” [9, 17, 20] by a German type foundry called URW is a well-known optical spacing software not for letter pairs but for text string. About font design software, the following link is useful: https://www.monotype.com/ resources/introduction-software-type-design.) It should be noted that heuristic rules are imperfect — after applying the rules, a final adjustment process by some kerning specialists is still required. For example, although FontForge is one of de-facto standard font design software tools, which have a heuristics-based automatic kerning mode, users need to revise its kerning result.
@@ 4.3 Comparative methods and evaluation metrics
  • “Average” determines the letter space between a certain letter pair (say, “AB”) by averaging the ground-truth spaces of that pair (“AB”) from all fonts in the training set. Consequently, the pair “AB” always has the same space regardless of font, whereas the spaces for “AB” and “IJ” are different. Although this is not a standard way for letter spacing, we introduce this to understand the importance of the letter space variations in different fonts.
@@ 4.5 Qualitative evaluation
These results, especially those with small MAEs (less than about two pixels), show that the set-wise model performs near-perfect space estimation. More importantly, the results are still acceptable even when larger MAEs. In fact, in Fig. 9, the word “UNIVERSITY” with the estimated spaces has no extra space to split the word into fragments and no under-estimated space to overlap adjacent letters, except for Mrs Sheppards Regular, where letters in a Handwriting style are overlapping. Another interesting result is Lekton Italic, which becomes the worst due to the under-estimated spaces of several pairs; however, the under-estimated spaces still seem also (or even more) natural.
@@ 6 Conclusion, Limitation, and Future Work
Our current limitation is the inflexibility of the estimation output. In this paper, we treated the default letter spaces as the ground-truth and tried to estimate the letter spaces similar to the ground-truth. In other words, we assumed that the appropriate letter space is uniquely determined for each letter pair of each font. However, typographic designers modify the letter spaces to create a particular impression. For example, letter spacing slightly wider than the default gives a more established and confident impression. In future work, such a space modification mechanism should be introduced into the current model to allow for more flexible kerning. We plan to apply our model to AI-generated fonts and expand our model to estimate letter spaces while considering three or more letters or even a paragraph (like the hz-program). In addition, we will modify our model to deal with font data in vector formats, although we used raster images. Finally, subjective evaluation by the kerning experts is mandatory.
Acknowledgment: This work was supported by JSPS KAKENHI Grant Number JP22H00540. We thank anonymous reviewers for their professional knowledge about kerning.